OpenStudy (anonymous):

Factorize fully 3(x+2)2-3x(x+2)

OpenStudy (luffingsails):

Need to get rid of parentheses by expanding and then recombining like terms: 3(x+2)2-3x(x+2) = (3x+6)2-3x^2-6x (3x+6)2-3x^2-6x = 6x+12-3x^2-6x 6x+12-3x^2-6x=12-3x^2 Does that make sense?

OpenStudy (anonymous):